# Visuomotor Policy Learning via Coarse-to-Fine Autoregressive Prediction

### v2

@@ 3.3 Implementation Details
Our primary focus is on the CARP algorithm, and we adopt a simple model architecture design.
Tokenization. Due to the disentangling of the action prediction, imprecise coarse-to-fine actions representation can decrease the upper limit of model performance. Considering the discontinuity of the most representation for rotation space like Euler angle or quaternion will increase the unstable training process, we utilize rotation6d [71] to get stable multi-scale tokens. And for the distance function $\operatorname{dist}(\cdot)$, we use cosine similarity rather than Euclidean distance which is the cause of unstable training based on our observations. In practice, due to the orthogonality between dimensions, we use a separate VQVAE [58] for each dimension of the action space, to gain stable training. All convolutions we used in CARP are 1D to capture the time-dimension features.
Autoregressive. We adopt the architecture of standard decoder-only transformers akin to GPT-2 [58, 9]. The state $\boldsymbol{s}_{O}$ is used to generate the initial coarse-scale token map and then is utilized as adaptive normalization [40] for the subsequent predictions. During training, we observe that incorporating Exponential Moving Average (EMA) [18] enhances both training stability and performance, yielding a 4-5% improvement, consistent with findings in [11]. During the inference, kv-caching can be used and no mask is needed, as shown in Fig. 10.
@@ 4 Experiment
In this section, we evaluate CARP on diverse robotics tasks, including state-based and image-based benchmarks in single-task and multi-task settings. CARP’s performance is assessed based on task success rates, inference speed, and model parameter scale. Additionally, we validate CARP’s practical effectiveness by deploying it on real-world tasks using a UR5e robotic arm, comparing its performance against state-of-the-art diffusion-based policies.
Our experiments are structured to address the following key research questions:
Figure 4: Single-Task Simulation Setup. We evaluate three tasks from the Robomimic [38] benchmark—Lift, Can, and Square—ordered by increasing difficulty, along with a Kitchen task [17] on the far left.
  • RQ1: Can CARP match the accuracy and robustness of current state-of-the-art diffusion-based policies?
  • RQ2: Does CARP maintain high inference efficiency and achieve fast convergence?
@@
4.1 Evaluation on Simulation Benchmark
We first evaluate CARP on a set of simulated tasks commonly used to benchmark diffusion-based models, assessing its ability to match their performance while significantly improving computational efficiency.
@@ 4.1 Evaluation on Simulation Benchmark
Table 2: Image-Based Simulation Results (Visual Policy). Results show that CARP consistently balances high performance and high efficiency. We highlight our results in light-blue.
Experimental Setup. We use the Robomimic [38] benchmark suite, which is widely adopted for evaluating diffusion-based models [11, 61, 42]. We evaluate CARP on both state-based and image-based datasets collected from expert demonstrations, with each task containing 200 demonstrations. The selected tasks are standard in single-task evaluations, as shown in Fig. 4 and Fig. 12. For evaluating the ability to learn multiple long-horizon tasks, we utilize the Franka Kitchen environment [17] which contains 7 objects for interaction and comes with a human demonstration dataset of 566 demonstrations, each completing 4 tasks in arbitrary order (see  Fig. 14 for details).
Baselines. We compare CARP with previous autoregressive policies as well as recent diffusion policies. Behavior Transformer (BET) [52] is an autoregressive policy with action discretization and correction mechanisms, similar to offset-based prediction. Implicit Behavior Cloning (IBC) [15] utilizes energy-based models for supervised robotic behavior learning. Diffusion Policy (DP) [11] combines a denoising process with action prediction, implemented in two variants: CNN-based (DP-C) and Transformer-based (DP-T).
@@
[…]
@@
Metrics. For each task, we evaluate the policies by running 50 trials with random initializations to compute the success rate and report the average success rate across the top 3 checkpoints. For the Kitchen task, success rates are reported with increasing difficulty, determined by the number of interactive objects (from p1 to p4). Inference speed is measured on an A100 GPU by averaging the time required for 400 action predictions (280 for the Kitchen task) over 5 runs to ensure robustness. We also record the parameter count for each model using the same PyTorch implementation interface.
@@
Figure 5: Visualization of the Trajectory and Refining Process. The left panel shows the final predicted trajectories for each task, with CARP producing smoother and more consistent paths than Diffusion Policy (DP). The right panel visualizes intermediate trajectories during the refinement process for CARP (top-right) and DP (bottom-right). DP displays considerable redundancy, resulting in slower processing and unstable training, as illustrated by 6 selected steps among 100 denoising steps. In contrast, CARP achieves efficient trajectory refinement across all 4 scales, with each step contributing meaningful updates.
Implementation Details. For baseline models, we follow the same implementation and training configurations provided by [11]. In state-based experiments, we set the observation horizon $O=2$ and the prediction horizon $H=16$ across all models. For image-based experiments, we set $O=1$ and $H=16$ for better transferability to real-world scenarios. As per the benchmark, only the first 8 actions in the prediction horizon are executed starting from the current step (see Appendix C). For CARP, we first train an action VQVAE model (see Sec. 3.1) following [29], using $V=512$, $C=8$, a batch size of 256, and 300 epochs per task Given a horizon $H=16$, we design multi-scale representations with scales of $1,2,3$ and $4$ to capture coarse-to-fine information across the action sequence. We then train an autoregressive GPT-2 style, decoder-only transformer (see Sec. 3.2), based on [58], using the same training settings as the benchmark, with a batch size of 256 for state-based experiments (4000 epochs) and a batch size of 64 for image-based experiments (3000 epochs).
@@
Results. As demonstrated in Tables 1, 2, and 3, CARP consistently achieves comparable performance to state-of-the-art diffusion models across both state-based and image-based tasks, answering RQ1. Notably, CARP significantly outperforms diffusion policies in terms of inference speed, being approximately 10 times faster, with only 1-5% of the parameters required by diffusion models, thereby strongly supporting RQ2 regarding CARP’s efficiency.
@@
Analysis. To further examine CARP’s stability and efficiency, we visualize spatial trajectories along the $xyz$ axes for the Can and Square tasks in Fig. 5. In each task’s left panel, CARP consistently reaches specific regions (light grey area) to perform task-related actions, such as positioning for object grasping or placement. Compared to diffusion-based models, CARP’s trajectories are smoother and more consistent, underscoring its stability and accuracy (supporting RQ1). The right panels of Fig. 5 compare CARP’s coarse-to-fine action predictions with 6 selected denoising steps of the diffusion model. CARP achieves accurate predictions within just 4 coarse-to-fine steps, whereas the diffusion model requires numerous denoising iterations, with many early steps introducing redundant computations. This analysis further supports RQ2 and highlights CARP’s efficiency and fast convergence advantage over diffusion policies in generating accurate actions with fewer refinement steps, as detailed in Sec. 3.2.
4.2 Evaluation on Multi-Task Benchmark
@@
We evaluate CARP on the MimicGen [39] multi-task simulation benchmark, widely used by the state-of-the-art Sparse Diffusion Policy (SDP) [61], to demonstrate CARP’s flexibility, a result of its GPT-style autoregressive design.
Figure 6: Multi-Task Simulation Setup. We evaluate eight tasks from the MimicGen [39] benchmark: Coffee, Hammer Cleanup, Mug Cleanup, Nut Assembly, Square, Stack, Stack Three, and Threading, listed left-to-right and top-to-bottom.
@@
Experimental Setup. MimicGen extends benchmark Robomimic [38] by including 1K–10K human demonstrations per task, with diverse initial state distributions for enhanced generalization in multi-task evaluation. It consists of 12 robosuite [72] tasks powered by MuJoCo and 4 high-precision tasks from Isaac Gym Factory. We select 8 robosuite tasks—Coffee, Hammer, Mug, Nut, Square, Stack, Stack Three, and Threading—for evaluation, each with 1K training trajectories, following the settings in SDP [61]. Visualization is shown in Fig. 6 and Fig. 13.
Baselines. We compare CARP with two baselines: Task-Conditioned Diffusion (TCD) [2, 34]: A basic diffusion-based multi-task policy and Sparse Diffusion Policy (SDP) [61]: A transformer-based diffusion policy that leverages Mixture of Experts (MoE) [53]. Both baselines are trained using visual inputs.
Metrics. Success rates are reported for each task as the average of the best three checkpoints. For Nut Assembly, partial success (e.g., placing one block inside the cylinder) is assigned a score of 0.5, while full success is scored as 1. For all other tasks, a score of 1 is awarded only when strict success criteria are fully satisfied. Additionally, we calculate the average success rate across all tasks. To evaluate model efficiency, we report both the parameter scale calculated by the Pytorch interface and the inference time required to predict 400 actions on a single A100 GPU.
Implementation Details. For CARP, we extend the single-task implementation by incorporating a task embedding as an additional condition, alongside the observation sequence $\boldsymbol{s}$. We also use a moderately deeper decoder-only transformer in GPT-2 style. CARP is trained with a batch size of 512 for 200 epochs on an A100 GPU. Baseline models follow the same training settings as SDP [61]. This minimal modification enables CARP to adapt to multi-task learning seamlessly.
Results. As shown in Tab. 4, CARP achieves up to a 25% average improvement in success rates compared to state-of-the-art diffusion-based policies, highlighting its strong performance. With minimal modification, CARP seamlessly transitions from single-task to multi-task learning, further demonstrating its flexibility, a benefit of its GPT-style architecture. Additionally, as shown in Tab. 4, CARP achieves over 10× faster inference speed and uses only 10% of the parameters compared to SDP. Leveraging its GPT-style autoregressive design [7, 47], CARP can seamlessly adapt to multi-task settings without complex design modifications [61], demonstrating the flexibility of this architecture, which strongly supports RQ3. These results strongly demonstrate CARP’s flexibility, solidifying its role as a high-performance and high-efficiency approach for visuomotor robotic policies.
@@ 4.3 Evaluation on Real-World
Figure 7: Real-World Setup. The left panel shows the environment used for the experiment and demonstration collection. The right panel shows the trajectory from the Cup and Bowl datasets.
Experimental Setup. To validate CARP’s real-world applicability, we design two manipulation tasks:
  1. 1.
    Cup: The robot must locate a cup on the table, pick it up, move to the right area of the table, and put it down steadily (top-right, Fig. 7).
  2. 2.
    Bowl: The robot needs to identify a smaller bowl and a larger pot on the table, pick up the bowl, and place it inside the pot (bottom-right, Fig. 7).
We use a UR5e robotic arm with a Robotiq-2f-85 gripper, equipped with two RGB cameras: one mounted on the wrist and one in a third-person perspective (left panel, Fig. 7). The robot is controlled through 6D end-effector positioning, with inverse kinematics for joint angle calculation. For teleoperation, we collected 50 human demonstration trajectories for each task using a 3D Connexion space mouse.
Baselines. As a baseline, we reproduce the CNN-based image-based diffusion policy [11], adapting the model’s input size to accommodate our observational setup.
Figure 8: Real-World Results (Visual Policy). We report the average success rate across 20 trials and the inference speed as action prediction frequency. CARP achieves competitive success rates with significantly faster inference compared to DP.
Metrics. For each trained policy, we report the average success rate across 20 trials per task, with the initial positions randomized. We also measure inference speed on an NVIDIA GeForce RTX 2060 GPU, reporting action prediction frequency in Hertz.
Implement Details. For both models, the input consists of current visual observations from the wrist and scene cameras (resolution: $120\times 160$), as well as proprioceptive data from the robotic arm. We execute 8 predicted actions out of a horizon of 16 predictions. We train the diffusion policy for 3000 epochs with a batch size of 64. For CARP, we use the same visual policy structure as in the simulation tasks, training the action VQVAE for 300 epochs with a batch size of 256, and the decoder-only transformer for 3000 epochs with a batch size of 64.
Results. As shown in Fig. 8 and Fig. 9, CARP achieves comparable or superior performance, with up to a 10% improvement in success rate over the diffusion policy across all real-world tasks, supporting RQ1. Additionally, CARP achieves approximately 8×8$\times$8 × faster inference than the baseline on limited computational resources, demonstrating its suitability for real-time robotic applications, thus supporting RQ2 (see Fig. 15 for further visualization).
Figure 9: Visualization of CAPR on Real-World Tasks. CARP generates smooth and successful trajectories for the Cup and Bowl tasks, with temporal progression from left to right.
@@ 5 Related Work
5.1 Visual Generation
Advancements in generative models for visual generation have significantly influenced the robotics community. Autoregressive models generate images in a raster-scan fashion using discrete tokens from image tokenizers [14, 20, 14, 60]. GPT-2-style transformers [29, 4, 48, 63] demonstrate strong performance by generating tokens sequentially. Recent work has scaled these models to billions of parameters, achieving impressive text-to-image synthesis results [69, 35] and robotic action generation [13, 54]. VAR [58] introduces a new next-scale autoregressive paradigm that shifts image representation from patches to scales. This framework [58], has been applied across tasks [33, 32, 67, 43, 66, 37, 16]. Studies [58, 56] show that autoregressive models can surpass diffusion models in achieving state-of-the-art performance, serving as a key inspiration for our work.
5.2 Visuomotor Policy Learning
Behavior cloning [8] has proven effective, especially in autonomous driving [41] and manipulation [68], offering a simpler alternative to complex reinforcement learning. Explicit policies directly map states or observations to actions [68], enabling efficient inference. However, they struggle with complex tasks. Techniques like action space discretization [65] and Mixture Density Networks (MDNs) [52, 12] have been proposed, though they face issues with exponential action space growth and hyperparameter sensitivity. Implicit policies, often based on Energy-Based Models (EBMs) [23, 15], offer flexibility but are challenging to train due to optimization instabilities. Diffusion models have been applied to robotic policy learning [11, 50, 2], showing their effectiveness for decision-making tasks. However, their multi-step denoising process is computationally expensive. Recent efforts focus on improving generalization [62], adapting to 3D environments [61, 64], and enhancing modularity via Mixture of Experts (MoE)  [61]. Consistency models [55, 36] accelerate inference but often compromise action prediction accuracy, along with inflexible model design.

### v3

@@ 3.3 Implementation Details
Here, we present the key implementation details of CARP.
Tokenization. Due to the disentangling of the action prediction, imprecise coarse-to-fine action representations can decrease the upper limit of model performance. Considering the discontinuity of the most representation for rotation space like Euler angle or quaternion will increase the unstable training process, we utilize rotation6d [73] to get stable multi-scale tokens. And for the distance function $\operatorname{dist}(\cdot)$, we use cosine similarity rather than Euclidean distance which is the cause of unstable training based on our observations. In practice, due to the orthogonality between dimensions, we use a separate VQVAE [60] for each dimension of the action space, to gain stable training. All convolutions we used in CARP are 1D to capture the time-dimension features.
Autoregressive. We adopt the architecture of standard decoder-only transformers akin to GPT-2 [60, 10]. The state $\boldsymbol{s}_{O}$ is used to generate the initial coarse-scale token map and then is utilized as adaptive normalization [42] for the subsequent predictions. During training, we observe that incorporating Exponential Moving Average (EMA) [19] enhances both training stability and performance, yielding a 4-5 % improvement, consistent with findings in [12]. During the inference, kv-caching can be used and no mask is needed (for further inference details, refer to Suppl. B).
@@ 4 Experiment
In this section, we evaluate CARP on diverse robotics tasks, including state-based and image-based benchmarks in single-task and multi-task settings. CARP’s performance is assessed in terms of task success rate, inference speed, and model parameter scale. Additionally, we validate CARP’s practical effectiveness by deploying it on real-world tasks using UR5e and Franka robotic arms, comparing its performance against state-of-the-art diffusion-based policies. Our experiments are structured to address the following key research questions (see Suppl. H for experimental implementation details):
  • RQ1: Can CARP achieve accuracy and robustness comparable to current state-of-the-art diffusion-based policies?
  • RQ2: Does CARP maintain high inference efficiency, characteristic of autoregressive models?
@@
Figure 4: Simulation Tasks Visualization. In the single-task setting, we evaluate Lift, Can, and Square, ordered by increasing difficulty, along with the Kitchen task for long-horizon evaluation and Push-T task for assessing multi-modal behavior. In the multi-task setting, we consider eight tasks: Coffee, Hammer Cleanup, Mug Cleanup, Nut Assembly, Square, Stack, Stack Three, and Threading, arranged from left to right and top to bottom.
4.1 Evaluation on Single-Task Simulations
We begin by evaluating CARP on a set of standard simulated tasks commonly used to benchmark diffusion-based policies, aiming to assess whether it can achieve comparable performance while significantly improving inference efficiency.
@@ 4.1 Evaluation on Single-Task Simulations
Table 2: Image-Based Simulation Results (Visual Policy). Results show that CARP consistently balances high performance and high inference efficiency. We highlight our results in light-blue.
[…]
Baselines. We compare CARP with previous autoregressive policies and recent diffusion policies. Behavior Transformer (BET) [54] is an autoregressive policy with action discretization and correction mechanisms, similar to offset-based prediction. We further extend BET with action chunking to improve performance, as shown in Fig. 2(a). Implicit Behavior Cloning (IBC) [16] utilizes energy-based models for supervised robotic behavior learning. Diffusion Policy (DP) [12] combines a denoising process with action prediction, implemented in two variants: CNN-based (DP-C) and Transformer-based (DP-T). Both follow the official implementation, employing DDPM with 100 denoising steps.
@@
Table 3: Multi-Stage Task Results (State Policy). In the Kitchen, pxx represents the success rate of interacting with $x$ or more objects. CARP outperforms BET, especially on challenging metrics like p4, and achieves competitive performance compared to DP, with fewer parameters and faster inference speed.
Figure 5: Multi-Modal Behavior Results (State Policy). On the Push-T task featuring multi-modal path options, CARP generates diverse, consistent predictions from the same initialization. Notably, it outperforms baselines in both task success and inference speed.
@@
Metrics. For each task, we evaluate policies by conducting 50 trials with random initializations, computing the success rate, and reporting the average over the top three checkpoints. In the Kitchen task, success rates are cumulative, requiring completion of previous levels to achieve the next (e.g., p2 requires completing p1 first). Inference speed is measured on an A100 GPU by averaging the time taken for predicting 400 actions (280 for Kitchen, 300 for Push-T) across five runs to ensure robustness. We also record each policy’s parameter count using the same PyTorch interface, excluding the visual encoder in the image-based setting.
@@
Figure 6: Visualization of the Trajectory and Refining Process. The left panel shows the final predicted trajectories for each task, with CARP producing smoother and more consistent paths than Diffusion Policy (DP). The right panel visualizes intermediate trajectories during the refinement process for CARP (top-right) and DP (bottom-right). DP displays considerable redundancy, resulting in slower processing and unstable action prediction, as illustrated by 6 selected steps among 100 denoising steps. In contrast, CARP achieves efficient trajectory refinement across all 4 scales, with each step contributing meaningful updates.
@@
Results. As demonstrated in Tables 1, 2, and 3, CARP consistently outperforms autoregressive policies and demonstrates comparable performance to diffusion policies across both state-based and image-based tasks, answering RQ1. Notably, CARP significantly outperforms diffusion policies in terms of inference speed, being approximately 10 times faster, with only 1-5% of the parameters required by diffusion models, thereby strongly supporting RQ2 regarding CARP’s efficiency. See Suppl. G for extra failure analysis. As shown in Fig. 5, CARP’s use of Cross-Entropy loss at each scale during training facilitates latent-space sampling at inference, enhancing multimodality and producing smoother actions. This leads to superior performance and faster inference, further supporting RQ1 and RQ2.
@@
Analysis. To further examine CARP’s stability and efficiency, we visualize spatial trajectories along the xyzxyz axes for the Can and Square tasks in Fig. 6. In each task’s left panel, CARP consistently reaches specific regions (light grey area) to perform task-related actions, such as positioning for object grasping or placement. Compared to diffusion-based policies, CARP’s trajectories are smoother and more consistent, underscoring its stability and accuracy (supporting RQ1). The right panels of Fig. 6 compare CARP’s coarse-to-fine action predictions with 6 selected denoising steps of the diffusion model. CARP achieves accurate predictions within just 4 coarse-to-fine steps, whereas the diffusion model requires numerous denoising iterations, with many early steps introducing redundant computations. This analysis further supports RQ2 and highlights CARP’s efficiency and fast inference convergence advantage over diffusion policies in generating accurate actions with fewer refinement steps, as detailed in Sec. 3.2. See Suppl. D, E, and F for more analysis.
4.2 Evaluation on Multi-Task Simulations
@@
We further evaluate CARP on the MimicGen [41] multi-task simulation benchmark, widely used by the state-of-the-art Sparse Diffusion Policy (SDP) [63], to demonstrate CARP’s flexibility, a result of its GPT-style autoregressive design.
@@
Experimental Setup. MimicGen extends benchmark Robomimic [40] by including 1K–10K human demonstrations per task, with diverse initial state distributions for enhanced generalization in multi-task evaluation. We select 8 robosuite [74] tasks for evaluation, each with 1K training trajectories, following the settings in SDP [63], as shown in Fig. 4. For CARP, we simply extend the single-task implementation by incorporating a task embedding as an additional condition alongside the observation sequence $\boldsymbol{s}$.
Baselines. We compare CARP with two baselines: task-conditioned diffusion (TCD) [2, 36]: A basic diffusion-based multi-task policy and sparse diffusion policy (SDP) [63]: A transformer-based diffusion policy that leverages mixture of experts (MoE) [55]. Both baselines are trained with visual inputs following their official implementation and settings.
Metrics. Success rates are reported for each task as the average of the best three checkpoints. For Nut Assembly, partial success (e.g., placing one block inside the cylinder) is assigned a score of 0.5, while full success is scored as 1. For all other tasks, a score of 1 is awarded only when strict success criteria are fully satisfied. Additionally, we calculate the average success rate across all tasks. To evaluate model efficiency, we report both parameter counts (excluding the identically configured visual encoder) and the inference time required to predict 400 actions on a single A100 GPU.
Results. As shown in Tab. 4, CARP achieves up to a 25% average improvement in success rates compared to state-of-the-art diffusion-based policies, highlighting its strong performance. Additionally, as shown in Tab. 4, CARP achieves over 10× faster inference speed and uses only 10% of the parameters compared to SDP. Leveraging its GPT-style autoregressive design [8, 49], CARP can seamlessly transition from single-task to multi-task settings with minimal structural modifications [63], demonstrating the flexibility of this architecture, which strongly supports RQ3. Considering fine-grained manipulation, such as Nut and Threading, CARP outperforms diffusion methods by up to 20%, demonstrating its strong applicability in these tasks (see Suppl. L for further analysis). These results strongly demonstrate CARP’s flexibility, solidifying its role as a high-performance and high-efficiency approach for visuomotor robotic policies.
@@ 4.3 Evaluation on Real-World
Experimental Setup. To validate CARP’s real-world applicability, we design two manipulation tasks: 1) Cup: The robot must locate a cup on the table, pick it up, move to the right area of the table, and put it down steadily. 2) Bowl: The robot needs to identify a smaller bowl and a larger pot on the table, pick up the bowl, and place it inside the pot. We use a UR5e robotic arm with a Robotiq-2f-85 gripper, equipped with two RGB cameras: one mounted on the wrist and one in a third-person perspective (left panel, Fig. 7). The robot is controlled through 6D end-effector positioning, with inverse kinematics for joint angle calculation. For teleoperation, we collected 50 human demonstration trajectories for each task using a 3D Connexion space mouse. We reproduce the image-based CNN-based Diffusion Policy as a baseline, as it has been shown to outperform current autoregressive policies, adapting the model’s input size to match our observational setup. CARP maintains the design with $K=4$, consistent with previous experiments (Suppl. J for ablation studies).
Metrics. For each trained policy, we report the average success rate across 20 trials per task, with the initial positions randomized. We also measure inference speed on an NVIDIA GeForce RTX 2060 GPU, reporting action prediction frequency in Hertz (Suppl. M for more visualization).
Results. As shown in the top-right table of Fig. 7, CARP achieves comparable or superior performance, with up to a 10% improvement in success rate over the Diffusion Policy across all real-world tasks, supporting RQ1. Additionally, CARP achieves approximately $8\times$ faster inference than the baseline on limited computational resources, demonstrating its suitability for real-time robotic applications, thus supporting RQ2 (Suppl. K for additional real-world experiments).
@@ 5 Related Work
Visual Generation. Advances in visual generative models have strongly influenced the robotics community. Autoregressive models generate images using discrete tokens from image tokenizers [22, 15, 62]. GPT-2-style transformers [31, 4, 50, 65] demonstrate strong performance by generating tokens sequentially. Recent work has scaled these models to achieve impressive text-to-image synthesis results [71, 37] and robotic action generation [14, 56]. VAR [60] introduces a new next-scale autoregressive paradigm that shifts image representation from patches to scales. This framework [60], has been applied across tasks [35, 34, 69, 45, 68, 39, 17]. Studies [60, 58] show that autoregressive models can surpass diffusion models in achieving compatible performance, serving as a key inspiration for the design of our approach.
Figure 7: Real-World Study. The left panel illustrates the environment used for both experimentation and demonstration collection. The bottom-right panel visualizes trajectories from the Cup and Bowl datasets. In the top-right panel, we present the average success rate over 20 trials alongside inference speed, measured as action prediction frequency. CARP achieves competitive success rates while significantly outperforming DP in inference speed.
Visuomotor Policy Learning. Behavior cloning [9] provides an effective approach, especially in autonomous driving [43] and manipulation [70], offering a simpler alternative to complex reinforcement learning. Explicit policies map observations to actions efficiently [70], but struggle with complex tasks. Solutions like action discretization [67] and mixture density networks [54, 13] mitigate this, yet suffer from action space explosion and hyperparameter sensitivity. Implicit policies, such as energy-based models [25, 16], offer greater flexibility but are harder to train due to optimization challenges. Diffusion models have proven effective for policy learning [12, 52, 2], but suffer from high computational cost due to multi-step denoising. Recent work aims to improve generalization [64], support 3D environments [63, 66], and enhance modularity via mixture of experts [63]. Consistency models [57, 38] accelerate inference but compromise action prediction accuracy, along with inflexible model design.